OpenStudy (anonymous):

Let f(x) = x2 + 6 and g(x) =x+8/x. Find ( g o f)(-7)

OpenStudy (anonymous):

@Directrix @Catlover5925

OpenStudy (catlover5925):

oh ok i got this, i would say if it is F of G plug in G first to get F and if it is G of F plug in G first then F

OpenStudy (anonymous):

im still lost lol

OpenStudy (catlover5925):

@Callisto @iGreen @sammixboo @AriPotta @uri @Jhannybean @confluxepic @Directrix @Diyadiya plz help explain this!!!!!!!

OpenStudy (igreen):

\((x^2 + 6)(\dfrac{x+8}{x})(-7)\)

OpenStudy (p0sitr0n):

g o f (-7) => g(f(-7)) so first find f(-7), then replace it in g() f(-7)=49+6=55 g(f(-7))=55+8/55=3033/55

OpenStudy (p0sitr0n):

assuming x2 means x squared

OpenStudy (anonymous):

mental abusive to humans lol

OpenStudy (anonymous):

so what do i do next after weget 3033/55

OpenStudy (anonymous):

@P0sitr0n

OpenStudy (p0sitr0n):

it is the answer

OpenStudy (p0sitr0n):

\[(g \; o \; f) (x) = g(f(x))\]

OpenStudy (anonymous):

it has to be a.63/55 b.295/49c.-57/7or d.384/7 @P0sitr0n

OpenStudy (anonymous):

@P0sitr0n

OpenStudy (anonymous):

@igreen help please

OpenStudy (catlover5925):